You intend to conduct an ANOVA with 4 groups in which each group will have the same number of subjects: n=16n=16. (This is referred to as a "balanced" single-factor ANOVA.)

What are the degrees of freedom for the numerator?
d.f.(treatment) =

What are the degrees of freedom for the denominator?
d.f.(error) =
Answer #1

degrees of freedom for the numerator =Number of groups (k)-1 =4-1 =3

degrees of freedom for the denominator =nk-n =16*4-4=60